AGw and GCS Commissioning Observing on 08/09 September 2008.

Observers: D. Miller, A. Rakich, N Cushing and J. Hill in Tucson. A gaggle of Lucifer people on the mountain
Telescope Operator: D. Huerta


Did not open due to clouds and high humidity

A new TCS build was installed and GCS Guiding and ActOpts were tested with dome closed. All went well.

The configuration file for the IRC Client was set to the system default on the LBTO account and the IRAF script were tested to confirm the correct configuration. Doug modified his IDL routines to read a separate (/home/LBTO/irtc_client.config) file which sets the Verbose mode to 2. The IDL routines were confirmed to work properly.

The IRS server on is now used for all irc requests.

Doug and Andrew re-crunched some field data to confirm the orientation and scale of the needed off-axis corrections to minimize the Wavefront Error on-axis when guiding and running the ActOpt loop on an off-axis star. A sign error was found in the IDL routines and corrected. Doug proceeded to compare the "correct" IDL corrections to those calculated in GCS. Not consistent, so Doug will work with Torsten tomorrow.

Nighttime Activities

Dome Closed all night

-- DougMiller - 10 Sep 2008
Topic revision: r1 - 10 Sep 2008, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback